OpenStudy (anonymous):

The functions f(x) and g(x) are described using the following equation and table: f(x) = -6(1.02)x x g(x) -1 -5 0 -3 1 -1 2 1 Which equation best compares the y-intercepts of f(x) and g(x)? The y-intercept of f(x) is equal to the y-intercept of g(x). The y-intercept of f(x) is equal to 2 times the y-intercept of g(x). The y-intercept of g(x) is equal to 2 times the y-intercept of f(x). The y-intercept of g(x) is equal to 2 plus the y-intercept of f(x). @iGreen. Last question then ill tell you what i get

OpenStudy (anonymous):

@Luigi0210 @ganeshie8

OpenStudy (anonymous):

@uri @camerondoherty

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

x=-1,0,1,2 g(x)=-5,-4,-3,-2,-1,1

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

https://www.desmos.com/calculator/nermtjb4ra

OpenStudy (anonymous):

Look at the y-intercepts of both graphs.

OpenStudy (anonymous):

What's the similarity?

OpenStudy (anonymous):

i dont see one but i think its B

OpenStudy (anonymous):

bc gx is going down

OpenStudy (anonymous):

Yep, you got it.

OpenStudy (anonymous):

y-intercept of g(x) = -3 y-intercept of f(x) = -6 -3 * 2 = -6 So f(x)'s y-intercept = 2 * g(x)'s y-intercept.

OpenStudy (anonymous):

okay well i got an 85 on my test

OpenStudy (anonymous):

c: